# Leptogenesis from magnetic helicity of gauged $\text{U}(1)_{B-L}$

### v1

@@ II.2 Evolution of the helicity
Next, we discuss the dynamics of the plasma with $B-L$ interaction and the $B-L$ gauge field after inflation, with instant reheating in mind. In this subsection, we focus on the case where the $B-L$ gauge field is massless. Throughout this paper, we require that the magnetohydrodynamics (MHD) description is valid for the $B-L$ gauge field and the magnetic helicity is conserved if the $B-L$ symmetry is not broken. We clarify the necessary conditions for these requirements in the following discussion. As is the case for the standard model hypercharge gauge field Giovannini and Shaposhnikov (1998b); Son (1999); Jedamzik et al. (1998); Banerjee and Jedamzik (2004); Domcke et al. (2019), the following MHD equation is to be satisfied;
@@
where $\vec{J}_{B-L}$ is the total $B-L$ current, $\sigma$ is the $B-L$ conductivity, $\vec{v}$ is the velocity of the fluid, $\eta_{m}athrm{vis}$ is the shear viscosity, $\rho$ is the energy density of the fluid, and $p$ is the pressure of the fluid. We have ignored the chiral magnetic effect Fukushima et al. (2008); Son and Surowka (2009); Neiman and Oz (2011) which does not change the evolution of the system unless the chiral plasma instability Joyce and Shaposhnikov (1997); Akamatsu and Yamamoto (2013); Rogachevskii et al. (2017); Schober et al. (2018); Kamada (2018) or the anomalous chirality cancellation Brandenburg et al. (2023a, b) becomes effective. In order to explain the present baryon asymmetry of the universe, the mechanism discussed in the subsequent subsections should take place. The $B-L$ magnetic field must be subdominant to the radiation energy density of the plasma, $B_{X}^{2}\ll\rho$. For simplicity, we assume all particles in the plasma are massless and in thermal equilibrium. Additionally, we assume that the typical scale of the dynamics, $L$, is much larger than the mean-free path of the $B-L$ charge carriers. The cosmic expansion does not appear in the MHD equations, because for massless gauge fields it can be removed by moving to the conformal frame Brandenburg et al. (1996). Here we use a conventional treatment following Ref. Domcke et al. (2019), which assumes that the typical spatial scales for the magnetic and velocity fields are the same. Although recent studies show that it is not always the case Uchida et al. (2023, 2024), this approach is enough to give a rough estimate.
Let us consider the validity of these equations. The first equation is the Ohm’s law for the $B-L$ current and can be qualitatively understood from the Drude model Arnold et al. (2000). In the Drude model, the conductivity of the $B-L$ current in the conformal frame is
@@
where $a(T)$ is the scale factor of the universe at $T$, $q_{i}$ is the $B-L$ charge of the charge carrier $i$, $T$ is the temperature of the universe, $g_{i}$ is the number of the internal degrees of freedom of $i$ and $\tau_{i}$ is the mean-free time for $i$. For the standard model particles, the dominant contribution comes from the right-handed charged leptons, where $\tau\sim 1/(g^{\prime 4}T)$, where $g^{\prime}$ is the $\text{U}(1)_{Y}$ gauge coupling constant. With new particles, such as a sterile neutrino, which interacts with the standard model particles only via the $B-L$ gauge field, $\tau$ can be longer and $\sigma$ can be bigger. However, as we have assumed $L\gg\tau$, there is an upper bound for $\sigma$, $\sigma\lesssim a(T)g_{B-L}^{2}T^{2}L.$ Thus, as a typical value, we take $\sigma\sim a(T)g_{B-L}^{2}T/g^{\prime 4}$ in the subsequent discussion.
The second equation is the Ampère’s law for the $B-L$ current; it is derived from one of the Maxwell’s equations,
×$\nabla\times\vec{B}_{X}=\vec{J}_{B-L}+\frac{\partial}{\partial t}\vec{E}_{X}.$ (II.25)
@@
Figure 1: The parameter space where the $B-L$ helicity is well-conserved, with the magnetic field created during the inflation. The vertical axis is the effective coupling constant of the $B-L$ gauge field, and the horizontal axis is the instability parameter $\xi$ between the inflaton and the $B-L$ gauge field. The black dashed lines are contours of the effective inflaton-$B-L$ instability parameter, $\xi_{\text{eff}}=3,4,\cdots 10$, from the left to the right. The blue, orange, and green shaded regions correspond to $H_{\mathrm{inf}}=10^{12},10^{10}$ and $10^{8}\,\text{GeV}$, respectively, where $H_{\mathrm{inf}}$ is the Hubble constant during the inflation. For the details of these regions, see the main text.
In Fig. 1, we illustrate the parameter space where the $B-L$ helicity is well-conserved, under the assumption that the magnetic field is generated during inflation and the reheating process is instantaneous. The light-shaded regions indicate where the energy density of the $B-L$ gauge field is sufficiently small (lower boundary) and where the helicity-to-entropy ratio is sufficiently large (upper boundary). Specifically, for the lower boundary, we require that the energy density of the $B-L$ gauge field is less than $1\%$ of the total energy density at the reheating temperature. For the upper boundary, we define the helicity-to-entropy ratio as
@@
where $s$ is the entropy density of the universe. We require that the helicity-to-entropy ratio is more than $\frac{79}{28}\eta_{0}$, where $\eta_{0}\simeq 8.7\times 10^{-11}$ is the observed baryon-to-entropy ratio in the current universe Aghanim et al. (2020), to ensure that the $B-L$ helicity decay would explain the present baryon asymmetry of the universe as we will discuss in the next subsection. The dark-shaded regions indicate where the magnetic Reynolds number is larger than the unity (Eq. (II.38)) such that the $B-L$ helicity is well-conserved among the corresponding light-shaded regions. In these regions, we show the region where the kinetic Reynolds number is larger than unity, $R_{e}>1$, as the meshed regions. We also show the contours of the effective inflaton-$B-L$ instability parameter, $\xi_{\text{eff}}=3,4,\cdots,10$, according to the equilibrium estimation reviewed in the previous subsection.
@@
The contours of $\xi_{\text{eff}}$ show that the effective instability parameter can be large if the gauge coupling $g_{B-L}$ is small. This is because the backreaction to the effective instability parameter is smaller for smaller $g_{B-L}$. For the larger $\xi_{\text{eff}}$, the helicity-to-entropy ratio can be larger despite the smaller gauge coupling. For the $B-L$ helicity to be well-conserved, the magnetic Reynolds number should be much larger than unity, $R_{m}\gg 1$. This requires that the magnetic field is sufficiently large, and therefore the viable parameter space lies in the smaller $g_{B-L}$ and larger $\xi_{\text{eff}}$ regions.
@@
Before moving to the next subsection, let us comment on the uncertainty of our discussion about the MHD evolution. Since the MHD equations describe a fully non-linear system, the estimates are not very precise to determine its evolution. While a sufficiently large magnetic Reynolds number is a necessary condition for the survival of the magnetic helicity, our estimate is based on rough order estimates and hence is not quantitatively precise. For example, the estimate of the velocity field at a small kinetic Reynolds number might not be correct since it is not clear if a turbulent dynamics takes place, which is essential for our order estimate. This is the reason why Ref. Domcke et al. (2023) adopted several ways to estimate the magnetic Reynolds number. Moreover, recent studies suggest that the typical length scale for the magnetic field and velocity field might be different Uchida et al. (2023, 2024), which violates the first assumption of our estimate. Therefore, one should regard the parameter space for the survival of the $B-L$ helicity in Fig. 1 as just a demonstration for a rough estimate. The origin of this uncertainty comes from the hierarchy between the magnetic and kinetic Reynolds number, $R_{m}/R_{e}\sim g^{2}_{B-L}/g^{\prime 8}$, for not so small $g_{B-L}$, which is difficult to be implemented in the numerical simulation. In order to determine precisely the condition for the helicity to be well-conserved, more precise studies in the case where the magnetic and kinetic Reynolds numbers are hierarchical are needed, but it is beyond the scope of the present study.
@@ II.3 Decay of the helicity
Let us discuss the fate of the $B-L$ helicity in the presence of a Higgs field with a negative mass squared. It is important to note that even if such a Higgs field exists, the phase of the system may differ from the one without helicity due to the anomaly equation, Eq. (II.3), which links the helicity to the fermion number Rubakov (1986). However, in our case of interest, where the helicity-to-entropy ratio is much smaller than unity, the phase of the system is not significantly affected by the helicity, and the $B-L$ symmetry is broken.
In the Higgs phase, the magnetic helicity can decay through two mechanisms: the imaginary part of its self-energy and Ohmic dissipation. The former would be analogous to the decay of a $B-L$ gauge boson particle in a vacuum. Consequently, we expect the decay rate to be similar to that of the $B-L$ gauge boson in a vacuum. The lifetime $\tau_{X}$ is thus given by
@@
To estimate the decay rate due to Ohmic dissipation, we first write down the equation of motion for the $B-L$ gauge field and the matter fields, which may replace Eq. (II.21) and Eq. (II.22) in the previous subsection. For the time being, let us consider the Minkowski spacetime, $a=1$ for simplicity. We will back to this point later. With the existence of a would-be Nambu-Goldstone boson $\phi_{M}$ for the $B-L$ gauge symmetry, the effective Lagrangian of the $B-L$ gauge field is
@@
where $m_{X}$ is the mass of the $B-L$ gauge field and $v_{B-L}$ is the VEV of the Higgs field breaking the $B-L$ gauge symmetry Here, for simplicity, we ignore the Chern-Simons term between $X_{\mu}$ and $\phi_{M}$.. The mass of the $B-L$ gauge field is $m_{X}=qg_{B-L}v_{B-L}$, where $q$ is the $B-L$ charge of the Higgs field. We may redefine the $B-L$ gauge field as $X_{\mu}\to X_{\mu}+g_{B-L}^{-1}v_{B-L}^{-1}\partial_{\mu}\phi_{M}$ to get rid of the would-be Nambu-Goldstone boson. The equation of motion of the $B-L$ gauge field is then
@@
×$\nabla\times\vec{B}_{X}-\frac{\partial}{\partial t}\vec{E}_{X}$ $=\vec{J}_{B-L}-m_{X}^{2}\vec{X}.$ (II.43)
@@
Here, again we have ignored the chiral magnetic effect. In this paper, we assume that the $B-L$ symmetry is Higgsed before the anomalous chirality cancellation Brandenburg et al. (2023a, b) would occur. The gauge transformation of the matter fields can absorb the redefinition, and the equations of motion of the matter fields are not changed. Thus, we assume that the Ohm’s law, Eq. (II.21), is still valid after the $B-L$ gauge symmetry is broken. The time dependence of the electric field is not negligible, and the electric conductivity may be different from the one before the $B-L$ gauge symmetry is broken. However, for the time being, we assume $\sigma\sim g_{B-L}^{2}T^{2}\tau\sim g_{B-L}^{2}T/g^{\prime 4}$ if the temperature $T$ is not much smaller than $v_{B-L}$. We will return to this point later.
@@
To further simplify the equation, we assume the typical spatial scale of the $B-L$ gauge field is much larger than the typical time scale of the $B-L$ gauge field, $L\gg T^{\prime}$. We will discuss the validity of this assumption later. Also, in the situation of our interest, the net $B-L$ charge is zero or, at least, much smaller than the current and we may neglect the term $X^{0}$. The equation then becomes
@@
Let us now make a few comments. First, we have assumed that the electric conductivity is not much different from the one before the $B-L$ gauge symmetry is broken. However, in the broken phase, the time dependence of the electric field can be much larger than the spatial dependence of the electric field; if the fluid is non-conducting, the time frequency of the electric field is $m_{X}$. When the gauge field is oscillating, the electric conductivity is different from the one for non-oscillating gauge field. According to the Drude model Ashcroft and Mermin (2011), the electric conductivity, $\sigma(\omega)$, for oscillating gauge field with the frequency $\omega$ is now complex and
@@
For $T\ll v_{B-L}$, $m_{X}\tau\gg 1$ and the fluid becomes non-conducting. Second, we have assumed that the typical spatial scale of the $B-L$ gauge field is much larger than the typical time scale of the $B-L$ gauge field. This is justified for
@@
We have so far ignored the cosmic expansion and the $B-L$ phase transition. Let us estimate when the helicity decays in the cosmic history and examine its validity. Here we assume that the gauge field obtains the mass via the Higgs mechanism,
@@
where $v_{B-L}(T)$ is the VEV of the Higgs field at the temperature $T$. We also assume that the $B-L$ phase transition is of the second order. $v_{B-L}(T)$ depends on the critical exponent of the $B-L$ phase transition, $\beta$, $v_{B-L}(T)\propto(T_{c}-T)^{\beta}$, where $T_{c}$ is the critical temperature of the $B-L$ phase transition. Let $T_{HD}$ be the temperature when the helicity decays. As we have discussed in this section, it is difficult to estimate the decay rate of the helicity precisely, but for simplicity, we assume that $T_{HD}$ is the temperature when the helicity decay time is equal to the Hubble time,
@@
Here, because $L\gtrsim H(T_{HD})^{-1}|\vec{v}|$ due to the MHD dynamics Banerjee and Jedamzik (2004)§§§Here we suppose that the coherence length of magnetic field is larger than the Alfvén scale. If we adopt the reconnection-driven turbulence Uchida et al. (2023, 2024), the coherence length can be shorter., the requirement, Eq. (II.50), is satisfied at least for $q\frac{v_{B-L}(T_{HD})}{T_{HD}}\sim\frac{1}{g^{\prime 2}}\sqrt{\frac{H(T_{HD})}{T_{HD}}}$. If we assume the Landau-Ginzburg theory,
@@
with $v_{B-L}^{0}\sim T_{c}$, we can conclude $T_{c}-T_{HD}\ll T_{c}$; the helicity quickly decays after the phase transition, in a shorter time scale than the Hubble time. Therefore, we expect that the effect of the cosmic expansion is negligible. If this condition is not satisfied, for example, if the mass of the $B-L$ gauge field is Stueckelberg-like The ’t Hooft anomaly may be canceled by the four-dimensional analogue of the Green-Schwarz mechanism Green and Schwarz (1984); Faddeev and Shatashvili (1986); Krasnikov (1985); Babelon et al. (1986); Harada and Tsutsui (1987); Preskill (1991), but the cutoff scale cannot be arbitrarily large in that case. and no phase transition occurs, the helicity may also decay due to cosmic expansion.

### v2

@@ II.2 Evolution of the helicity
Next, we discuss the dynamics of the plasma with BLB-L interaction and the BLB-L gauge field after inflation, with instant reheating in mind. In this subsection, we focus on the case where the BLB-L gauge field is massless. Throughout this paper, we require that the magnetohydrodynamics (MHD) description is valid for the BLB-L gauge field and the magnetic helicity is conserved if the BLB-L symmetry is not broken. We clarify the necessary conditions for these requirements in the following discussion. As is the case for the standard model hypercharge gauge field Giovannini and Shaposhnikov (1998b); Son (1999); Jedamzik et al. (1998); Banerjee and Jedamzik (2004); Domcke et al. (2019), the following MHD equation is to be satisfied;
@@
where $\vec{J}_{B-L}$ is the total BLB-L current, $\sigma$ is the BLB-L conductivity, $\vec{v}$ is the velocity of the fluid, $\eta_{\text{vis}}$ is the shear viscosity, $\rho$ is the energy density of the fluid, and $p$ is the pressure of the fluid. We have ignored the chiral magnetic effect Fukushima et al. (2008); Son and Surowka (2009); Neiman and Oz (2011) which does not change the evolution of the system unless the chiral plasma instability Joyce and Shaposhnikov (1997); Akamatsu and Yamamoto (2013); Rogachevskii et al. (2017); Schober et al. (2018); Kamada (2018) or the anomalous chirality cancellation Brandenburg et al. (2023a, b) becomes effective. In order to explain the present baryon asymmetry of the universe, the mechanism discussed in the subsequent subsections should take place. The BLB-L magnetic field must be subdominant to the radiation energy density of the plasma, $B_{X}^{2}\ll\rho$. For simplicity, we assume all particles in the plasma are massless and in thermal equilibrium. Additionally, we assume that the typical scale of the dynamics, $L$, is much larger than the mean-free path of the BLB-L charge carriers. The cosmic expansion does not appear in the MHD equations, because for massless gauge fields it can be removed by moving to the conformal frame Brandenburg et al. (1996). Here we use a conventional treatment following Ref. Domcke et al. (2019), which assumes that the typical spatial scales for the magnetic and velocity fields are the same. Although recent studies show that it is not always the case Uchida et al. (2023, 2024), this approach is enough to give a rough estimate.
Let us consider the validity of these equations. The first equation is the Ohm’s law for the BLB-L current and can be qualitatively understood from the Drude model Arnold et al. (2000). In the Drude model, the conductivity of the BLB-L current in the conformal frame is
@@
where $a(T)$ is the scale factor of the universe at $T$, $q_{i}$ is the BLB-L charge of the charge carrier $i$, $T$ is the temperature of the universe, $g_{i}$ is the number of the internal degrees of freedom of $i$ and $\tau_{i}$ is the mean-free time for $i$. For the standard model particles, the dominant contribution comes from the right-handed charged leptons, where $\tau\sim 1/(g^{\prime 4}T)$, where $g^{\prime}$ is the $\text{U}(1)_{Y}$ gauge coupling constant. With new particles, such as a sterile neutrino, which interacts with the standard model particles only via the BLB-L gauge field, $\tau$ can be longer and $\sigma$ can be bigger. However, as we have assumed $L\gg\tau$, there is an upper bound for $\sigma$, $\sigma\lesssim a(T)g_{B-L}^{2}T^{2}L.$ Thus, as a typical value, we take $\sigma\sim a(T)g_{B-L}^{2}T/g^{\prime 4}$ in the subsequent discussion.
The second equation is the Ampère’s law for the BLB-L current; it is derived from one of the Maxwell’s equations,
$\nabla\times\vec{B}_{X}=\vec{J}_{B-L}+\frac{\partial}{\partial t}\vec{E}_{X}.$ (II.25)
@@
Figure 1: The parameter space where the BLB-L helicity is well-conserved, with the magnetic field created during the inflation. The vertical axis is the effective coupling constant of the BLB-L gauge field, and the horizontal axis is the instability parameter $\xi$ between the inflaton and the BLB-L gauge field. The black dashed lines are contours of the effective inflaton-BLB-L instability parameter, $\xi_{\text{eff}}=3,4,\cdots 10$, from the left to the right. The blue, orange, and green shaded regions correspond to $H_{\mathrm{inf}}=10^{12},10^{10}$ and $10^{8}\,\text{GeV}$, respectively, where $H_{\mathrm{inf}}$ is the Hubble constant during the inflation. For the details of these regions, see the main text.
In Fig. 1, we illustrate the parameter space where the BLB-L helicity is well-conserved, under the assumption that the magnetic field is generated during inflation and the reheating process is instantaneous. The light-shaded regions indicate where the energy density of the BLB-L gauge field is sufficiently small (lower boundary) and where the helicity-to-entropy ratio is sufficiently large (upper boundary). Specifically, for the lower boundary, we require that the energy density of the BLB-L gauge field is less than $1\%$ of the total energy density at the reheating temperature. For the upper boundary, we define the helicity-to-entropy ratio as
@@
where $s$ is the entropy density of the universe. We require that the helicity-to-entropy ratio is more than $\frac{79}{28}\eta_{0}$, where $\eta_{0}\simeq 8.7\times 10^{-11}$ is the observed baryon-to-entropy ratio in the current universe Aghanim et al. (2020), to ensure that the BLB-L helicity decay would explain the present baryon asymmetry of the universe as we will discuss in the next subsection. The dark-shaded regions indicate where the magnetic Reynolds number is larger than the unity (Eq. (II.38)) such that the BLB-L helicity is well-conserved among the corresponding light-shaded regions. In these regions, we show the region where the kinetic Reynolds number is larger than unity, $R_{e}>1$, as the meshed regions. We also show the contours of the effective inflaton-BLB-L instability parameter, $\xi_{\text{eff}}=3,4,\cdots,10$, according to the equilibrium estimation reviewed in the previous subsection.
@@
The contours of $\xi_{\text{eff}}$ show that the effective instability parameter can be large if the gauge coupling $g_{B-L}$ is small. This is because the backreaction to the effective instability parameter is smaller for smaller $g_{B-L}$. For the larger $\xi_{\text{eff}}$, the helicity-to-entropy ratio can be larger despite the smaller gauge coupling. For the BLB-L helicity to be well-conserved, the magnetic Reynolds number should be much larger than unity, $R_{m}\gg 1$. This requires that the magnetic field is sufficiently large, and therefore the viable parameter space lies in the smaller $g_{B-L}$ and larger $\xi_{\text{eff}}$ regions.
@@
Before moving to the next subsection, let us comment on the uncertainty of our discussion about the MHD evolution. Since the MHD equations describe a fully non-linear system, the estimates are not very precise to determine its evolution. While a sufficiently large magnetic Reynolds number is a necessary condition for the survival of the magnetic helicity, our estimate is based on rough order estimates and hence is not quantitatively precise. For example, the estimate of the velocity field at a small kinetic Reynolds number might not be correct since it is not clear if a turbulent dynamics takes place, which is essential for our order estimate. This is the reason why Ref. Domcke et al. (2023) adopted several ways to estimate the magnetic Reynolds number. Moreover, recent studies suggest that the typical length scale for the magnetic field and velocity field might be different Uchida et al. (2023, 2024), which violates the first assumption of our estimate. Therefore, one should regard the parameter space for the survival of the BLB-L helicity in Fig. 1 as just a demonstration for a rough estimate. The origin of this uncertainty comes from the hierarchy between the magnetic and kinetic Reynolds number, $R_{m}/R_{e}\sim g^{2}_{B-L}/g^{\prime 8}$, for not so small $g_{B-L}$, which is difficult to be implemented in the numerical simulation. In order to determine precisely the condition for the helicity to be well-conserved, more precise studies in the case where the magnetic and kinetic Reynolds numbers are hierarchical are needed, but it is beyond the scope of the present study.
@@ II.3 Decay of the helicity
Let us discuss the fate of the BLB-L helicity in the presence of a Higgs field with a negative mass squared. It is important to note that even if such a Higgs field exists, the phase of the system may differ from the one without helicity due to the anomaly equation, Eq. (II.3), which links the helicity to the fermion number Rubakov (1986). However, in our case of interest, where the helicity-to-entropy ratio is much smaller than unity, the phase of the system is not significantly affected by the helicity, and the BLB-L symmetry is broken.
In the Higgs phase, the magnetic helicity can decay through two mechanisms: the imaginary part of its self-energy and Ohmic dissipation. The former would be analogous to the decay of a BLB-L gauge boson particle in a vacuum. Consequently, we expect the decay rate to be similar to that of the BLB-L gauge boson in a vacuum. The lifetime $\tau_{X}$ is thus given by
@@
To estimate the decay rate due to Ohmic dissipation, we first write down the equation of motion for the BLB-L gauge field and the matter fields, which may replace Eq. (II.21) and Eq. (II.22) in the previous subsection. For the time being, let us consider the Minkowski spacetime, $a=1$ for simplicity. We will back to this point later. With the existence of a would-be Nambu-Goldstone boson $\phi_{M}$ for the BLB-L gauge symmetry, the effective Lagrangian of the BLB-L gauge field is
@@
where $m_{X}$ is the mass of the BLB-L gauge field and $v_{B-L}$ is the VEV of the Higgs field breaking the BLB-L gauge symmetry §§§Here, for simplicity, we ignore the Chern-Simons term between $X_{\mu}$ and $\phi_{M}$.. The mass of the BLB-L gauge field is $m_{X}=qg_{B-L}v_{B-L}$, where $q$ is the BLB-L charge of the Higgs field. We may redefine the BLB-L gauge field as $X_{\mu}\to X_{\mu}+g_{B-L}^{-1}v_{B-L}^{-1}\partial_{\mu}\phi_{M}$ to get rid of the would-be Nambu-Goldstone boson. The equation of motion of the BLB-L gauge field is then
@@
$\nabla\times\vec{B}_{X}-\frac{\partial}{\partial t}\vec{E}_{X}$ $=\vec{J}_{B-L}-m_{X}^{2}\vec{X}.$ (II.43)
@@
Here, again we have ignored the chiral magnetic effect. In this paper, we assume that the BLB-L symmetry is Higgsed before the anomalous chirality cancellation Brandenburg et al. (2023a, b) would occur. The gauge transformation of the matter fields can absorb the redefinition, and the equations of motion of the matter fields are not changed. Thus, we assume that the Ohm’s law, Eq. (II.21), is still valid after the BLB-L gauge symmetry is broken. The time dependence of the electric field is not negligible, and the electric conductivity may be different from the one before the BLB-L gauge symmetry is broken. However, for the time being, we assume $\sigma\sim g_{B-L}^{2}T^{2}\tau\sim g_{B-L}^{2}T/g^{\prime 4}$ if the temperature $T$ is not much smaller than $v_{B-L}$. We will return to this point later.
@@
To further simplify the equation, we assume the typical spatial scale of the BLB-L gauge field is much larger than the typical time scale of the BLB-L gauge field, $L\gg T^{\prime}$. We will discuss the validity of this assumption later. Also, in the situation of our interest, the net BLB-L charge is zero or, at least, much smaller than the current and we may neglect the term $X^{0}$. The equation then becomes
@@
Let us now make a few comments. First, we have assumed that the electric conductivity is not much different from the one before the BLB-L gauge symmetry is broken. However, in the broken phase, the time dependence of the electric field can be much larger than the spatial dependence of the electric field; if the fluid is non-conducting, the time frequency of the electric field is $m_{X}$. When the gauge field is oscillating, the electric conductivity is different from the one for non-oscillating gauge field. According to the Drude model Ashcroft and Mermin (2011), the electric conductivity, $\sigma(\omega)$, for oscillating gauge field with the frequency $\omega$ is now complex and
@@
For $T\ll v_{B-L}$, $m_{X}\tau\gg 1$ and the fluid becomes non-conducting. Second, we have assumed that the typical spatial scale of the BLB-L gauge field is much larger than the typical time scale of the BLB-L gauge field. This is justified for
@@
We have so far ignored the cosmic expansion and the BLB-L phase transition. Let us estimate when the helicity decays in the cosmic history and examine its validity. Here we assume that the gauge field obtains the mass via the Higgs mechanism,
@@
where $v_{B-L}(T)$ is the VEV of the Higgs field at the temperature $T$. We also assume that the BLB-L phase transition is of the second order. $v_{B-L}(T)$ depends on the critical exponent of the BLB-L phase transition, $\beta$, $v_{B-L}(T)\propto(T_{c}-T)^{\beta}$, where $T_{c}$ is the critical temperature of the BLB-L phase transition. Let $T_{HD}$ be the temperature when the helicity decays. As we have discussed in this section, it is difficult to estimate the decay rate of the helicity precisely, but for simplicity, we assume that $T_{HD}$ is the temperature when the helicity decay time is equal to the Hubble time,
@@
Here, because $L\gtrsim H(T_{HD})^{-1}|\vec{v}|$ due to the MHD dynamics Banerjee and Jedamzik (2004)Here we suppose that the coherence length of magnetic field is larger than the Alfvén scale. If we adopt the reconnection-driven turbulence Uchida et al. (2023, 2024), the coherence length can be shorter., the requirement, Eq. (II.50), is satisfied at least for $q\frac{v_{B-L}(T_{HD})}{T_{HD}}\sim\frac{1}{g^{\prime 2}}\sqrt{\frac{H(T_{HD})}{T_{HD}}}$. If we assume the Landau-Ginzburg theory,
@@
with $v_{B-L}^{0}\sim T_{c}$, we can conclude $T_{c}-T_{HD}\ll T_{c}$; the helicity quickly decays after the phase transition, in a shorter time scale than the Hubble time. Therefore, we expect that the effect of the cosmic expansion is negligible. If this condition is not satisfied, for example, if the mass of the BLB-L gauge field is Stueckelberg-like The ’t Hooft anomaly may be canceled by the four-dimensional analogue of the Green-Schwarz mechanism Green and Schwarz (1984); Faddeev and Shatashvili (1986); Krasnikov (1985); Babelon et al. (1986); Harada and Tsutsui (1987); Preskill (1991), but the cutoff scale cannot be arbitrarily large in that case. and no phase transition occurs, the helicity may also decay due to cosmic expansion.